OpenStudy (anonymous):

Solve this system using substitution and give your answer as a point: y = 3x + 2 and y = 4x - 7. In other words, specify (x,y). Don't include any spaces.

OpenStudy (anonymous):

4x-7=3x+2 -3x -3x x-7=2 +7 +7 x=9 y=4(9)-7=29 Therefore; (9,29)